Brenda’s Bar and Grill has current liabilities of $27 million. Cash makes up 10 percent of the current assets and accounts receivable makes up another 40 percent of current assets. Brenda’s current ratio is 2.1 times. Calculate the value of inventory listed on the firm’s balance sheet. (Enter your answer in millions of dollars rounded to 2 decimal places.)
Answer #1

Current ratio=current assets/current liabilities

current assets=(27*2.1)=$56.7million

Cash=(56.7*10%)=$5.67 million

AR=(56.7*40%)=$22.68 million

current assets=Cash+AR+inventory

inventory=(56.7-5.67-22.68) million

=$28.35 million.